# Robust Methods for Multiscale Coarse Approximations of Diffusion Models in Perforated Domains

### v3

@@ Abstract
Parameter Notation
Parameter Definition
$\left(D_{j}\right)_{j=1,\ldots,N}$
Nonoverlapping partitioning of polygonal domain $D$
$\Omega_{S}$
Union of all perforations
$\left(\Omega_{j}\right)_{j=1,\ldots,N}$
Nonoverlapping partitioning of $\Omega=D\setminus\overline{\Omega_{S}}$
$\left(\Omega_{j}^{\prime}\right)_{j=1,\ldots,N}$
Overlapping partitioning such that $\Omega_{j}\subset\Omega_{j}^{\prime}$.
$\Gamma$
Coarse skeleton $\bigcup_{j=1,\ldots,N}\partial\Omega_{j}\setminus\partial\Omega_{S}$
$\left(e_{k}\right)_{k=1,\ldots,N_{e}}$
Nonoverlapping partitioning of $\Gamma$ into coarse edges
${\cal V}$
Set of coarse grid nodes $\bigcup_{k=1,\ldots,N_{e}}\partial e_{k}$
$H$
Max length over all coarse edges, $\max_{k=1,\ldots,N_{e}}|e_{k}|$.
$\mathcal{H}_{j}$
Maximum (length or width) distance between minimal and maximal coordinates of $\Omega_{j}$
$N_{\Omega}$
Set of internal triangulation nodes
${\bf A}$
$N_{\Omega}\times N_{\Omega}$ matrix derived from finite element triangulation of Poisson equation
${\bf R}_{H}$
Discrete coarse matrix of harmonic basis functions
${\bf R}_{j}^{\prime}$
Boolean restriction matrices for $\Omega_{j}^{\prime}$
${\bf R}_{j}^{\prime}$
Boolean restriction matrices for $\Omega_{j}^{\prime}$
$\overline{{\bf R}}_{j}$
Boolean restriction matrices for the closure of $\Omega_{j}$, given by $\overline{\Omega}_{j}$
@@ 1 Introduction and Model Problem
Numerical modeling of overland flows plays an increasingly important role in predicting, anticipating, and controlling floods. Anticipating these flood events can aid in the positioning of protective systems including dams, dikes, or rainwater drainage networks. One of the challenges of this numerical modeling of urban floods is that small structural features (buildings, walls, etc.) may significantly affect the water flow. In particular, the representation of the structural features in the numerical model impacts both the timing and extent of the flood Yu and Lane (2006a, b); Abily et al. (2016).
Modern terrain survey techniques including photogrammetry and Laser Imaging, Detection, and Ranging (LIDAR) allow us to acquire high-resolution topographic data for urban areas. The data set used in this article has been provided by Métropole Nice Côte d’Azur (MNCA) and allows for an infra-metric description of the urban geometries Andres (2012). From the hydraulic perspective, these structural features can be assumed to be essentially impervious, and therefore represented as perforations (holes) in the model domain.
c In this article, we consider a linear diffusion model. Let $D$ be an open simply connected polygonal domain in $\mathbb{R}^{2}$. We denote by $\left(\Omega_{S,k}\right)_{k}$ a finite family of perforations in $D$ such that each $\Omega_{S,k}$ is an open connected polygonal subdomain of $D$. The perforations are mutually disjoint, that is $\overline{\Omega_{S,k}}\cap\overline{\Omega_{S,l}}=\emptyset$ for any $k\neq l$. We denote $\Omega_{S}=\bigcup_{k}\Omega_{S,k}$ and $\Omega=D\setminus\overline{\Omega_{S}}$, assuming that the family $\left(\Omega_{S,k}\right)_{k}$ is such that $\Omega$ is connected. Note that the latter assumption implies that $\Omega_{S,k}$ are simply connected. We are interested in the boundary value problem given by
@@
In the context of urban flood modeling, $u$ represents the flow potential (pressure head) and $\left(\Omega_{S,k}\right)_{k}$ can be thought of as a family of impervious structures such as buildings, walls, or other similar structures. Although the linear model (1) is overly simplified to be directly used for urban flood assessment, the more general nonlinear elliptic or parabolic models are common in free surface flow simulations. Such models arise from Shallow Water equations either by neglecting the inertia terms Alonso et al. (2008) or within the context of semi-implicit Froude-robust time discretizations Casulli (1990). Depending on the geometrical complexity of the computational domain, the numerical resolution of (1) may become increasingly challenging. The typical model domain (see for example Figure 1) resulting from the realistic description of the urban environment will contain numerous perforations that are represented on different scales. The resolution of these structures can result in extremely small computational elements. Thus, we wish to employ numerical strategies that consider two levels of space discretization (see Figure 2 for a visual of both levels). Specifically, we employ a multiscale method to solve (1) on these urban geometries, aiming to achieve computational efficiency compared to classical fine-scale solution methods. Like other multiscale methods, this also allows for parallel implementation to further improve efficiency.
@@
There exists extensive literature on problems involving highly oscillatory coefficients or multiscale geometrical features, including for example the Localized Orthogonal Decomposition (LOD) method Målqvist and Peterseim (2014) and Heterogeneous Multiscale methods (HMM) Weinan and Engquist (2003). Additionally, to solve these heterogeneous problems, there exist multiscale partition of unity methods such as the Generalized Finite Element Method (GFEM) Babuska et al. (1994) and the Multiscale Spectral Generalized Finite Element Method (Ms-GFEM) Babuska and Lipton (2011); Ma et al. (2022). In Babuska and Lipton (2011); Ma et al. (2022), the optimal approximation spaces are constructed locally by solving generalized eigenvalue problems over a set of overlapping subdomains. The partition of unity functions are used to glue the local contributions into a globally conforming coarse approximation space. We note that these methods vary from the approach presented here as we consider nonoverlapping coarse cells and do not rely on eigenproblems.
@@
Since the perforations in our model problem are structured, we can exploit the structural features of the heterogeneities. The method we choose to consider here is closer to the Multiscale Hybrid-Mixed Method (MHM) Araya et al. (2013), the Multiscale Finite Element Method (MsFEM) Hou and Wu (1997), or polytopal methods such as Virtual Element Methods (VEM) Brezzi et al. (2014). In MsFEM, the domain is partitioned into coarse cells, allowing the method to take advantage of parallel computing. Basis functions are then computed numerically based on the PDE on each coarse cell; in this way, our methodology is similar to MsFEM. In comparison to the classical MsFEM, our method leads to a larger coarse space. Compared to VEM, the major difference is that we numerically compute the approximation of the locally harmonic basis. By doing so, we manage to incorporate singular functions (corresponding to the corners of the domain) into the coarse space. We are also able to deal with very general polygonal cells (not star-shaped, not simply connected, etc.), which differs from VEM. Additionally, we willingly avoid using a method of fundamental solutions of any kind, because of our long-term motivation in problems more complex than the linear model problem (1). The MHM method combines ideas from the Mixed Finite Element Method (MFEM) and MsFEM. Compared to MHM, where Neumann traces are approximated by piecewise polynomials, our method is formulated in terms of approximate Dirichlet traces. We remark as well the existence of Generalized Multiscale Finite Element Methods (GMsFEM) Efendiev et al. (2013), which were created to generalize MsFEM methods by constructing coarse trial spaces via a spectral decomposition of snapshot spaces.
The second contribution of the article is the combination of the proposed coarse approximation with local subdomain solves in a two-level domain decomposition (DD) method. The idea of DD methods is to partition the domain into multiple subdomains and solve on each subdomain locally, gluing the local solutions together at the end to obtain a global solution. This results in an efficient iterative solver for the algebraic system resulting from the fine-scale finite element method. The obtained algorithm can also be thought as a way of improving the precision of the original coarse approximation in the spirit of iterative multiscale methods (see e.g. Hajibeygi et al. (2008)). We note that furthermore, the discrete Trefftz coarse space can also be combined with domain decomposition methods to provide a two-level preconditioner for Krylov methods. The coarse component provides robustness of the preconditioned Krylov method in terms of iteration counts with respect to the number of subdomains. It also provides an additional acceleration when compared to the iterative method. Specifically on perforated domains with small and numerous perforations, the authors of Le Bris et al. (2014) introduced an MsFEM method for diffusion problems with Dirichlet boundary conditions imposed on the perforation and domain boundaries, with an error estimate provided. In Degond et al. (2015); Bris et al. (2019), this method was extended to advection-diffusion problems, with Bris et al. (2019) imposing both Dirichlet and Neumann boundary conditions on the perforation boundaries. In Le Bris et al. (2014); Bris et al. (2019); Degond et al. (2015), Crouzeix-Raviart type boundary conditions were imposed on the local problems to provide robustness with respect to the position of the perforations. As well, the addition of bubble functions is included. The authors of Brown and Taralova (2016) also provides an MsFEM method for perforated domains, providing numerical results and analysis for problems posed on domains with numerous small, regular perforations. In Brown and Taralova (2016), Neumann boundary conditions are posed on the perforation boundaries, with the assumption that the coarse grid does not intersect with the perforations along the edges for analysis purposes (although this assumption is not necessary numerically). Aside from classical MsFEM methods, the authors of Henning and Ohlberger (2009) introduced a HMM method to solve elliptic homogenization problems in perforated domains, with periodic perforations required for analysis purposes. Additionally, the authors of Chung et al. (2016a, b); Spiridonov et al. (2019) proposed a GMsFEM method for perforated domains for small, regular perforations. While literature on DD approaches to solve model problems in perforated domains is limited, the model problem can be thought of as the extreme limit case of the elliptic model containing highly contrasting coefficients with zero conductivity on the perforations. Two-level domain decomposition methods have been extensively studied for such heterogeneous problems. There are many classical results for coarse spaces that are constructed to resolve the jumps of the coefficients; see Dryja et al. (1994, 1996); Mandel and Brezina (1996) for further details. Robust coarse spaces have constructed using the ideas from MsFEM in Aarnes and Hou (2002); Graham et al. (2007). Approaches to obtain a robust coarse space without careful partitioning of the subdomains include spectral coarse spaces such as those given in Galvis and Efendiev (2010); Nataf et al. (2011); Dolean et al. (2012); Spillane et al. (2014). The authors of Scheichl et al. (2012) discuss these varied coefficient problems in the case where the coarse grid is not properly aligned with the heterogeneities. The combination of spectral and MsFEM methods can be found in Gander et al. (2015), where the authors enrich the MsFEM coarse space with eigenfunctions along the edges. Additionally, the family of GDSW (Generalized Dryja, Smith, Widlund) methods Dohrmann et al. (2008) employ energy-minimizing coarse spaces and can be used to solve heterogeneous problems on less regular domains. These spaces are discrete in nature and involve both edge and nodal basis functions. To deal with coefficient jumps in highly heterogeneous problems, an adaptive GDSW coarse space was introduced in Heinlein et al. (2018). The present article expands upon two published conference papers by the same authors Boutilier et al. (2022, 2023). Compared to the previous works, we consider a higher order Trefftz space and provide a detailed proof for the error estimate. In addition, we provide extended numerical experiments regarding the convergence of a two-level domain decomposition method based on the Trefftz coarse space. The article is laid out as follows. In Section 2, we introduce the continuous Trefftz coarse space and discuss its approximation properties. In Section 3, we introduce the Trefftz coarse space in its discrete matrix form. With this, we provide the matrix forms of the coarse approximation and domain decomposition methods. The two-level domain decomposition methods are presented as both an iterative solver and a preconditioner for Krylov methods. In Section 4, we provide numerical results for the Trefftz space used as a coarse approximation, in an iterative Schwarz method, and as a preconditioner for Krylov methods. We provide numerical results for two different types of model domains; the first domain is a simplified domain with one perforation at the corner, and the second type of model domain is a realistic urban domain with numerous perforations of various shape. Section 5 concludes with a summary and brief description of future work.
@@ 2.1 Coarse Mesh and Space Decomposition
By definition, $H^{1}_{\Delta}(\Omega)$ is orthogonal to $H^{1}_{\Gamma}(\Omega)$. Since $H^{1}_{\Gamma}(\Omega)$ is a closed subspace of $H^{1}(\Omega)$, we deduce that $H^{1}(\Omega)=H^{1}_{\Delta}(\Omega)\oplus H^{1}_{\Gamma}(\Omega)$ (see e.g. Rudin et al. (1973)). In other words, a given function $v\in H^{1}(\Omega)$ admits a unique decomposition into $v_{\Delta}+v_{b}$, where $v_{\Delta}\in H_{\Delta}^{1}(\Omega)$, $v_{b}\in H^{1}_{\Gamma}(\Omega)$ and $(v_{\Delta},v_{b})_{H^{1}(\Omega)}=0$. Although for simplicity we will call $v_{\Delta}$ the “locally harmonic” or “piece-wise” harmonic component of $v$, we wish to stress that the space $H^{1}_{\Delta}(\Omega)$ also contains information about the normal traces of $v_{\Delta}$ over $\partial\Omega_{S}$. The function $v_{b}$ will be referred to as the local or “bubble” component of $v$.
@@ 2.3 Error Analysis
See Proposition 1.5 and 1.12 of Ern et al. (2004). See also Lemma 4.2 and Remark 4 of Beir ao da Veiga et al. (2016) for estimates with explicit dependency on $p$ and higher dimension. ∎
@@
See Theorem 1 of Brezis and Mironescu (2018) for details. ∎
@@
The proof is quite basic, for the full proof we refer e.g. to Lemma 5 of Brenner et al. (2016). ∎

### v4

@@ Abstract
Parameter Notation Parameter Definition
$\left(D_{j}\right)_{j=1,\ldots,N}$ Nonoverlapping partitioning of polygonal domain $D$
$\Omega_{S}$ Union of all perforations
$\left(\Omega_{j}\right)_{j=1,\ldots,N}$ Nonoverlapping partitioning of $\Omega=D\setminus\overline{\Omega_{S}}$
$\left(\Omega_{j}^{\prime}\right)_{j=1,\ldots,N}$ Overlapping partitioning such that $\Omega_{j}\subset\Omega_{j}^{\prime}$.
$\Gamma$ Coarse skeleton $\bigcup_{j=1,\ldots,N}\partial\Omega_{j}\setminus\partial\Omega_{S}$
$\left(e_{k}\right)_{k=1,\ldots,N_{e}}$ Nonoverlapping partitioning of $\Gamma$ into coarse edges
${\cal V}$ Set of coarse grid nodes $\bigcup_{k=1,\ldots,N_{e}}\partial e_{k}$
$H$ Max length over all coarse edges, $\max_{k=1,\ldots,N_{e}}|e_{k}|$.
$\mathcal{H}_{j}$ Maximum (length or width) distance between minimal and maximal coordinates of $\Omega_{j}$
$N_{\Omega}$ Set of internal triangulation nodes
${\bf A}$ $N_{\Omega}\times N_{\Omega}$ matrix derived from finite element triangulation of Poisson equation
${\bf R}_{H}$ Discrete coarse matrix of harmonic basis functions
${\bf R}_{j}^{\prime}$ Boolean restriction matrices for $\Omega_{j}^{\prime}$
${\bf R}_{j}^{\prime}$ Boolean restriction matrices for $\Omega_{j}^{\prime}$
$\overline{{\bf R}}_{j}$ Boolean restriction matrices for the closure of $\Omega_{j}$, given by $\overline{\Omega}_{j}$
@@ 1 Introduction and Model Problem
Numerical modeling of overland flows plays an increasingly important role in predicting, anticipating, and controlling floods. Anticipating these flood events can aid in the positioning of protective systems including dams, dikes, or rainwater drainage networks. One of the challenges of this numerical modeling of urban floods is that small structural features (buildings, walls, etc.) may significantly affect the water flow. In particular, the representation of the structural features in the numerical model impacts both the timing and extent of the flood [1, 2, 3].
Modern terrain survey techniques including photogrammetry and Laser Imaging, Detection, and Ranging (LIDAR) allow us to acquire high-resolution topographic data for urban areas. The data set used in this article has been provided by Métropole Nice Côte d’Azur (MNCA) and allows for an infra-metric description of the urban geometries [4]. From the hydraulic perspective, these structural features can be assumed to be essentially impervious, and therefore represented as perforations (holes) in the model domain.
In this article, we consider a linear diffusion model. Let $D$ be an open simply connected polygonal domain in $\mathbb{R}^{2}$. We denote by $\left(\Omega_{S,k}\right)_{k}$ a finite family of perforations in $D$ such that each $\Omega_{S,k}$ is an open connected polygonal subdomain of $D$. The perforations are mutually disjoint, that is $\overline{\Omega_{S,k}}\cap\overline{\Omega_{S,l}}=\emptyset$ for any $k\neq l$. We denote $\Omega_{S}=\bigcup_{k}\Omega_{S,k}$ and $\Omega=D\setminus\overline{\Omega_{S}}$, assuming that the family $\left(\Omega_{S,k}\right)_{k}$ is such that $\Omega$ is connected. Note that the latter assumption implies that $\Omega_{S,k}$ are simply connected.
We are interested in the boundary value problem given by
@@
In the context of urban flood modeling, $u$ represents the flow potential (pressure head) and $\left(\Omega_{S,k}\right)_{k}$ can be thought of as a family of impervious structures such as buildings, walls, or other similar structures. Although the linear model (1) is overly simplified to be directly used for urban flood assessment, the more general nonlinear elliptic or parabolic models are common in free surface flow simulations. Such models arise from Shallow Water equations either by neglecting the inertia terms [5] or within the context of semi-implicit Froude-robust time discretizations [6].
Depending on the geometrical complexity of the computational domain, the numerical resolution of (1) may become increasingly challenging. The typical model domain (see for example Figure 1) resulting from the realistic description of the urban environment will contain numerous perforations that are represented on different scales. The resolution of these structures can result in extremely small computational elements. Thus, we wish to employ numerical strategies that consider two levels of space discretization (see Figure 2 for a visual of both levels). Specifically, we employ a multiscale method to solve (1) on these urban geometries, aiming to achieve computational efficiency compared to classical fine-scale solution methods. Like other multiscale methods, this also allows for parallel implementation to further improve efficiency.
@@
There exists extensive literature on problems involving highly oscillatory coefficients or multiscale geometrical features, including for example the Localized Orthogonal Decomposition (LOD) method [7] and Heterogeneous Multiscale methods (HMM) [8]. Additionally, to solve these heterogeneous problems, there exist multiscale partition of unity methods such as the Generalized Finite Element Method (GFEM) [9] and the Multiscale Spectral Generalized Finite Element Method (Ms-GFEM) [10, 11]. In [10, 11], the optimal approximation spaces are constructed locally by solving generalized eigenvalue problems over a set of overlapping subdomains. The partition of unity functions are used to glue the local contributions into a globally conforming coarse approximation space. We note that these methods vary from the approach presented here as we consider nonoverlapping coarse cells and do not rely on eigenproblems.
@@
Since the perforations in our model problem are structured, we can exploit the structural features of the heterogeneities. The method we choose to consider here is closer to the Multiscale Hybrid-Mixed Method (MHM) [12], the Multiscale Finite Element Method (MsFEM) [13], or polytopal methods such as Virtual Element Methods (VEM) [14]. In MsFEM, the domain is partitioned into coarse cells, allowing the method to take advantage of parallel computing. Basis functions are then computed numerically based on the PDE on each coarse cell; in this way, our methodology is similar to MsFEM. In comparison to the classical MsFEM, our method leads to a larger coarse space. Compared to VEM, the major difference is that we numerically compute the approximation of the locally harmonic basis. By doing so, we manage to incorporate singular functions (corresponding to the corners of the domain) into the coarse space. We are also able to deal with very general polygonal cells (not star-shaped, not simply connected, etc.), which differs from VEM. Additionally, we willingly avoid using a method of fundamental solutions of any kind, because of our long-term motivation in problems more complex than the linear model problem (1). The MHM method combines ideas from the Mixed Finite Element Method (MFEM) and MsFEM. Compared to MHM, where Neumann traces are approximated by piecewise polynomials, our method is formulated in terms of approximate Dirichlet traces. We remark as well the existence of Generalized Multiscale Finite Element Methods (GMsFEM) [15], which were created to generalize MsFEM methods by constructing coarse trial spaces via a spectral decomposition of snapshot spaces.
The second contribution of the article is the combination of the proposed coarse approximation with local subdomain solves in a two-level domain decomposition (DD) method. The idea of DD methods is to partition the domain into multiple subdomains and solve on each subdomain locally, gluing the local solutions together at the end to obtain a global solution. This results in an efficient iterative solver for the algebraic system resulting from the fine-scale finite element method. The obtained algorithm can also be thought as a way of improving the precision of the original coarse approximation in the spirit of iterative multiscale methods (see e.g. [16]). We note that furthermore, the discrete Trefftz coarse space can also be combined with domain decomposition methods to provide a two-level preconditioner for Krylov methods. The coarse component provides robustness of the preconditioned Krylov method in terms of iteration counts with respect to the number of subdomains. It also provides an additional acceleration when compared to the iterative method.
Specifically on perforated domains with small and numerous perforations, the authors of [17] introduced an MsFEM method for diffusion problems with Dirichlet boundary conditions imposed on the perforation and domain boundaries, with an error estimate provided. In [18, 19], this method was extended to advection-diffusion problems, with [19] imposing both Dirichlet and Neumann boundary conditions on the perforation boundaries. In [17, 19, 18], Crouzeix-Raviart type boundary conditions were imposed on the local problems to provide robustness with respect to the position of the perforations. As well, the addition of bubble functions is included. The authors of [20] also provides an MsFEM method for perforated domains, providing numerical results and analysis for problems posed on domains with numerous small, regular perforations. In [20], Neumann boundary conditions are posed on the perforation boundaries, with the assumption that the coarse grid does not intersect with the perforations along the edges for analysis purposes (although this assumption is not necessary numerically). Aside from classical MsFEM methods, the authors of [21] introduced a HMM method to solve elliptic homogenization problems in perforated domains, with periodic perforations required for analysis purposes. Additionally, the authors of [22, 23, 24] proposed a GMsFEM method for perforated domains for small, regular perforations.
While literature on DD approaches to solve model problems in perforated domains is limited, the model problem can be thought of as the extreme limit case of the elliptic model containing highly contrasting coefficients with zero conductivity on the perforations. Two-level domain decomposition methods have been extensively studied for such heterogeneous problems. There are many classical results for coarse spaces that are constructed to resolve the jumps of the coefficients; see [25, 26, 27] for further details. Robust coarse spaces have constructed using the ideas from MsFEM in [28, 29]. Approaches to obtain a robust coarse space without careful partitioning of the subdomains include spectral coarse spaces such as those given in [30, 31, 32, 33]. The authors of [34] discuss these varied coefficient problems in the case where the coarse grid is not properly aligned with the heterogeneities. The combination of spectral and MsFEM methods can be found in [35], where the authors enrich the MsFEM coarse space with eigenfunctions along the edges. Additionally, the family of GDSW (Generalized Dryja, Smith, Widlund) methods [36] employ energy-minimizing coarse spaces and can be used to solve heterogeneous problems on less regular domains. These spaces are discrete in nature and involve both edge and nodal basis functions. To deal with coefficient jumps in highly heterogeneous problems, an adaptive GDSW coarse space was introduced in [37].
The present article expands upon two published conference papers by the same authors [38, 39]. Compared to the previous works, we consider a higher order Trefftz space and provide a detailed proof for the error estimate. In addition, we provide extended numerical experiments regarding the convergence of a two-level domain decomposition method based on the Trefftz coarse space. The article is laid out as follows. In Section 2, we introduce the continuous Trefftz coarse space and discuss its approximation properties. In Section 3, we introduce the Trefftz coarse space in its discrete matrix form. With this, we provide the matrix forms of the coarse approximation and domain decomposition methods. The two-level domain decomposition methods are presented as both an iterative solver and a preconditioner for Krylov methods. In Section 4, we provide numerical results for the Trefftz space used as a coarse approximation, in an iterative Schwarz method, and as a preconditioner for Krylov methods. We provide numerical results for two different types of model domains; the first domain is a simplified domain with one perforation at the corner, and the second type of model domain is a realistic urban domain with numerous perforations of various shape. Section 5 concludes with a summary and brief description of future work.
@@ 2.1 Coarse Mesh and Space Decomposition
By definition, $H^{1}_{\Delta}(\Omega)$ is orthogonal to $H^{1}_{\Gamma}(\Omega)$. Since $H^{1}_{\Gamma}(\Omega)$ is a closed subspace of $H^{1}(\Omega)$, we deduce that $H^{1}(\Omega)=H^{1}_{\Delta}(\Omega)\oplus H^{1}_{\Gamma}(\Omega)$ (see e.g. [40]). In other words, a given function $v\in H^{1}(\Omega)$ admits a unique decomposition into $v_{\Delta}+v_{b}$, where $v_{\Delta}\in H_{\Delta}^{1}(\Omega)$, $v_{b}\in H^{1}_{\Gamma}(\Omega)$ and $(v_{\Delta},v_{b})_{H^{1}(\Omega)}=0$. Although for simplicity we will call $v_{\Delta}$ the “locally harmonic” or “piece-wise” harmonic component of $v$, we wish to stress that the space $H^{1}_{\Delta}(\Omega)$ also contains information about the normal traces of $v_{\Delta}$ over $\partial\Omega_{S}$. The function $v_{b}$ will be referred to as the local or “bubble” component of $v$.
@@ 2.3 Error Analysis
See Proposition 1.5 and 1.12 of [41]. See also Lemma 4.2 and Remark 4 of [42] for estimates with explicit dependency on $p$ and higher dimension. ∎
@@
See Theorem 1 of [43] for details. ∎
@@
The proof is quite basic, for the full proof we refer e.g. to Lemma 5 of [44]. ∎